# ELFATT: Efficient Linear Fast Attention for Vision Transformers

### v1

@@ IV-D Ablation Study of Number of Levels Using the ELFATT Module
To validate the effect of the number of ELFATT modules used to replace the vanilla softmax-based attention (VaniATT) modules at different levels of vision transformers, we compared the performance of CSWin-T24181 using a different number of ELFATT modules at different levels. As shown in Table IV, with the number of levels using ELFATT modules increasing, the inference speed increases without using FlashAttention-2 acceleration. The difference between CSWin-T24181 using 3 levels of ELFATT modules and 4 levels of ELFATT modules is not significant. The fourth level of CSWin-T24181 is composed of only one module and the sequence length is only 49 which is too short to achieve a significant acceleration effect. Another thing that can be found in Table IV is that with an increase in the number of levels using ELFATT modules, CSWin-T24181 using FlashAttention-2 acceleration achieves the fastest inference speed when using two levels of ELFATT modules. Because the sequence lengths of the first two levels are 3136 and 784, respectively, which are significantly longer than the last two levels (the $3^{\rm rd}$ level: 196, and the $4^{\rm th}$ level: 49). That is also the reason why some efficient attention mechanisms, such as Agent and FLatten, only replace some of levels by their efficient attention modules. We also observed that with an increasing number of levels using ELFATT modules, the performance shows a gradual decline. To address this defect, we introduced a hybrid architecture in the third level which replaces half of the ELFATT modules at this level by vanilla softmax-based attention modules. The order of composition is as follows: VaniATT-ELFATT-...-VaniATT-ELFATT. The reason is that the third level of the vision transformer is usually much deeper than other levels and the sequence of this level is also much shorter than the first two levels. VaniATT at this level will not affect the speed too much and can help the model to converge faster. Swin-T-ELFATT and CSWin-B36292-ELFATT also use a pure ELFATT architecture in the first two levels and a hybrid architecture in the third level. All variants of CSWin-T24181 using ELFATT modules to replace VaniATT modules are faster than CSWin-T24181-GLOBAL of which all levels are composed of VaniATT modules.
TABLE VI: The comparison of object detection performance of all methods on MS COCO 2017. Note: FLOPs are calculated using an input size of $1280\times 800$. ‘1$\times$’ denotes the fine-tuning training schedule with 12 epochs and ‘3$\times$×MS’ represents fine-tuning using the multiscale training schedule with 36 epochs. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. APb denotes box average precision and APm denotes mask average precision. Inference throughput is obtained using a batch size of 1 with mixed precision on a single NVIDIA H20 (96 GB) GPU. “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Mask-RCNN [48] 1$\times$ schedule
[…]
CSWin-T24181-Agent 46.8 68.9 51.3 42.3 65.9 45.3 19/20 imgs/s 40M 273.87G/254.51G
CSWin-T24181-EFFATT 46.1 68.3 50.5 41.9 65.5 45.3 32/36 imgs/s 40M 329.95G/255.20G
CSWin-T24181-ELFATT 47.0 69.2 51.4 42.6 66.4 45.9 30/33 imgs/s 40M 334.86G/254.40G
CSWin-T24181-FLatten 46.6 68.8 51.0 42.2 65.7 45.3 26/28 imgs/s 41M 274.41G/255.05G
CSWin-T24181-GLOBAL 47.0 69.1 51.9 42.6 66.1 45.9 7/18 imgs/s 40M 1712.76G/253.56G
CSWin-T24181-LOCAL 46.5 68.5 51.0 42.1 65.6 45.3 26/28 imgs/s 40M 281.45G/253.57G
Swin-T-Agent 44.6 67.5 48.7 40.7 64.4 43.4 5/— imgs/s 48M 278.42G/ —
Swin-T-EFFATT 44.7 67.0 48.9 41.1 64.0 44.4 40/46 imgs/s 48M 301.89G/261.95G
Swin-T-ELFATT 46.1 68.3 50.8 42.1 65.4 45.3 39/45 imgs/s 50M 311.39G/266.43G
Swin-T-FLatten 44.2 67.3 48.5 40.2 63.8 43.0 41/— imgs/s 49M 266.43G/ —
Swin-T-GLOBAL 45.4 67.9 49.7 41.6 65.0 44.8 6/17 imgs/s 48M 2106.75G/260.48G
Swin-T-LOCAL 42.7 65.2 46.8 39.3 62.2 42.2 45/— imgs/s 48M 267.01G/ —
ConvNeXt-T 44.2 66.6 48.3 40.1 63.3 42.8 44/— imgs/s 48M 262.29G/ —
VMamba-T 47.4 69.5 52.0 42.7 66.3 46.0 35/— imgs/s 50M 271.16G/ —
Mask-RCNN [48] 3$\times$×MS schedule
[…]
CSWin-T24181-Agent 49.3 70.8 53.9 43.9 67.9 47.3 19/20 imgs/s 40M 273.87G/254.51G
CSWin-T24181-EFFATT 48.5 70.0 53.2 43.4 67.3 46.9 32/36 imgs/s 40M 329.95G/255.20G
CSWin-T24181-ELFATT 49.4 70.9 54.4 44.0 68.0 47.5 30/33 imgs/s 40M 334.86G/254.40G
CSWin-T24181-FLatten 48.9 70.8 53.5 43.9 67.9 47.3 26/28 imgs/s 41M 274.41G/255.05G
CSWin-T24181-GLOBAL 48.8 70.0 53.5 43.6 67.4 47.1 7/18 imgs/s 40M 1712.76G/253.56G
CSWin-T24181-LOCAL 49.3 70.8 54.3 44.0 67.8 47.5 26/28 imgs/s 40M 281.45G/253.57G
Swin-T-Agent 47.3 69.5 51.9 42.7 66.4 46.2 5/— imgs/s 48M 278.42G/ —
Swin-T-EFFATT 47.6 69.4 52.6 42.7 65.9 46.1 40/46 imgs/s 48M 301.89G/261.95G
Swin-T-ELFATT 48.5 70.4 53.4 43.6 67.3 47.3 39/45 imgs/s 50M 311.39G/266.43G
Swin-T-FLatten 46.5 68.5 50.8 42.1 65.4 45.1 41/— imgs/s 49M 266.43G/ —
Swin-T-GLOBAL 48.0 70.0 52.7 43.3 67.0 46.8 6/17 imgs/s 48M 2106.75G/260.48G
Swin-T-LOCAL 46.0 68.1 50.3 41.6 65.1 44.9 45/— imgs/s 48M 267.01G/ —
ConvNeXt-T 46.2 67.9 50.8 41.7 65.0 44.9 44/— imgs/s 48M 262.29G/ —
VMamba-T 48.9 70.6 53.6 43.7 67.7 46.8 35/— imgs/s 50M 271.16G/ —
IV-E Ablation Study of Different Combinations of $c_{1}$ and $c_{2}$
@@
Table V shows the performance comparison of CSWin-T24181-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. When $c_{1}=1\times c$ and $c_{2}=0\times c$, ELFATT becomes EFFATT, and when $c_{1}=0\times c$ and $c_{2}=1\times c$, ELFATT becomes a local window-based attention mechanism which can be regarded as a simpler version of the local window-based attention mechanism used in Swin. As shown in Table V, $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ achieve better performance than other combinations of $c_{1}$ and $c_{2}$. Without using FlashAttention-2 acceleration, the difference between $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ in terms of speed is not significant. Using FlashAttention-2 acceleration, $c_{1}=0\times c$ and $c_{2}=1\times c$, and $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$ achieve close speed and are significantly faster than other combinations.
TABLE VII: The comparison of semantic segmentation performance of all methods on ADE20K. Note: The best values are in bold, ‘mAcc’ denotes mean class accuracy, and ‘mIoU’ denotes mean intersection over union. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. FLOPs are calculated using an input size of $512\times 2048$. ‘160k’ denotes the 160k fine-tuning iterations. Inference throughput is obtained using a batch size of 1 with mixed precision on a single NVIDIA H20 (96 GB) GPU. “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
UperNet [49] 160k
Method mAcc (%) mIoU (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-T24181-Agent 60.8 48.5 16/17 imgs/s 50M 953.60G/929.64G
CSWin-T24181-EFFATT 60.6 48.8 28/33 imgs/s 50M 1008.73G/930.34G
CSWin-T24181-ELFATT 61.2 49.6 28/32 imgs/s 50M 1014.26G/929.53G
CSWin-T24181-FLatten 61.4 49.3 25/27 imgs/s 51M 954.15G/930.19G
CSWin-T24181-GLOBAL 61.1 48.8  6/14 imgs/s 50M 2458.75G/928.67G
CSWin-T24181-LOCAL 61.1 49.6 26/28 imgs/s 50M 963.38G/928.68G
Swin-T-Agent 58.5 46.7 4/— imgs/s 61M 957.50G/ —
Swin-T-EFFATT 58.3 46.7 35/39 imgs/s 60M 981.22G/939.35G
Swin-T-ELFATT 59.3 47.7 34/38 imgs/s 62M 991.27G/943.94G
Swin-T-FLatten 57.0 44.8 35/— imgs/s 61M 944.62G/ —
Swin-T-GLOBAL 59.3 47.8 5/14 imgs/s 60M 2873.79G/937.84G
Swin-T-LOCAL 55.6 44.5 38/— imgs/s 60M 945.66G/ —
ConvNeXt-T 58.3 46.1 37/— imgs/s 60M 939.69G/ —
VMamba-T 59.3 47.9 34/— imgs/s 62M 948.78G/ —
IV-F Object Detection Performance
Table VI shows the object detection performance of all methods on MS COCO 2017. Using the Mask-RCNN [48] 1$\times$ schedule, the vanilla softmax-based attention mechanism and ELFATT significantly outperform other attention mechanisms. ELFATT exhibits slightly better performance than the vanilla softmax-based attention mechanism when using CSWin-T24181 as the backbone and significantly outperforms the vanilla softmax-based attention mechanism when using Swin-T as the backbone. Under the backbone of CSWin-T24181, ELFATT offers a 4.3x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 acceleration and is still 1.8x faster than the vanilla softmax-based attention mechanism using FlashAttention-2 acceleration. When using Swin-T as the backbone, ELFATT offers a 6.5x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 acceleration and is still 2.6x faster than the vanilla softmax-based attention mechanism using FlashAttention-2 acceleration. Using the Mask-RCNN [48] 3$\times$ multiscale training schedule, ELFATT still achieves the best performance. The vanilla softmax-based attention mechanism still outperforms most linear attention mechanisms when using Swin-T as the backbone. In the 1$\times$ schedule, ELFATT using CSWin-T24181 as the backbone achieves close performance compared to VMamba-T and significantly outperforms ConvNeXt-T. In the 3$\times$ multiscale training schedule, ELFATT using CSWin-T24181 as the backbone outperforms ConvNeXt-T and VMamba-T in terms of object detection performance.
TABLE VIII: The comparison of inference throughput of different attention mechanisms obtained on ImageNet-1K (Platform: NVIDIA Jetson AGX Orin; Batch Size: 128; Mode: 60W (Orin MAXN)).
Method Res. FPS (FP32) Speedup ratio FPS (Mixed precision) Speedup ratio
CSWin-T24181-Agent $224^{2}$ 158 imgs/s 1.5x 295 imgs/s 1.3x
CSWin-T24181-EFFATT $224^{2}$ 156 imgs/s 1.4x 305 imgs/s 1.3x
CSWin-T24181-ELFATT $224^{2}$ 174 imgs/s 1.6x 355 imgs/s 1.6x
CSWin-T24181-FLatten $224^{2}$ 138 imgs/s 1.3x 229 imgs/s 1.0x
CSWin-T24181-GLOBAL $224^{2}$ 108 imgs/s 1.0x 298 imgs/s 1.3x
CSWin-T24181-LOCAL $224^{2}$ 167 imgs/s 1.5x 309 imgs/s 1.3x
EfficientViT-B2 $288^{2}$ 198 imgs/s 1.8x 282 imgs/s 1.2x
TABLE IX: The comparison of inference throughput of different attention mechanisms obtained on ImageNet-1K (Platform: NVIDIA Jetson Nano; Batch Size: 1; Mode: 10W (Nano MAXN)).
Method Res. FPS (FP32) Speedup ratio FPS (Mixed precision) Speedup ratio
CSWin-T24181-Agent $224^{2}$ 4.9 imgs/s 1.5x 4.5 imgs/s 1.3x
CSWin-T24181-EFFATT $224^{2}$ 5.8 imgs/s 1.8x 5.5 imgs/s 1.6x
CSWin-T24181-ELFATT $224^{2}$ 6.2 imgs/s 2.0x 5.8 imgs/s 1.7x
CSWin-T24181-FLatten $224^{2}$ 4.7 imgs/s 1.5x 4.0 imgs/s 1.1x
CSWin-T24181-GLOBAL $224^{2}$ 3.2 imgs/s 1.0x 3.5 imgs/s 1.0x
CSWin-T24181-LOCAL $224^{2}$ 5.0 imgs/s 1.6x 4.2 imgs/s 1.2x
EfficientViT-B2 $288^{2}$ 5.5 imgs/s 1.7x 5.7 imgs/s 1.6x
Figure 4: The speed comparison of different attention mechanisms across various power modes from 5W to 60W by using NVIDIA Jetson Nano or NVIDIA Jetson AGX Orin.
IV-G Semantic Segmentation Performance
Table VII shows the comparison of semantic segmentation performance of all methods on ADE20K. The vanilla softmax-based attention mechanism and ELFATT achieve close mean class accuracy (mAcc) and mean intersection over union (mIoU), and significantly outperform other attention mechanisms when using Swin-T as the backbone. ELFATT, FLatten, the vanilla softmax-based attention mechanism, and the local window-based attention mechanism [23] achieve close mean class accuracy when using CSWin-T24181 as the backbone. However, ELFATT and the local window-based attention mechanism [23] significantly outperform other attention mechanisms in terms of mean intersection over union. Without using FlashAttention-2 acceleration, ELFATT offers a nearly 5x speedup over the vanilla softmax-based attention mechanism using CSWin-T24181 as the backbone and a 7x speedup over the vanilla softmax-based attention mechanism using Swin-T as the backbone. Even using FlashAttention-2 acceleration, under the CSWin-T24181 backbone, the vanilla softmax-based attention mechanism is still 2x slower than ELFATT without using FlashAttention-2 acceleration and 2.3x slower than ELFATT using FlashAttention-2 acceleration. Under the Swin-T backbone, the vanilla softmax-based attention mechanism is still 2.4x slower than ELFATT without using FlashAttention-2 acceleration and 2.7x slower than ELFATT using FlashAttention-2 acceleration. The speed of ELFATT is almost the same as the speed of EFFATT which is a real linear attention mechanism. However, ELFATT achieves significantly higher mean class accuracy and mean intersection over union than those of EFFATT. ELFATT using CSWin-T24181 as the backbone outperforms ConvNeXt-T and VMamba-T in terms of semantic segmentation performance.
IV-H Speed Comparison on Edge GPUs
@@
In addition to high-performance computing applications, model deployment should also be evaluated in emerging fields such as robotic vision, unmanned aerial vehicles (UAV), and autonomous driving. These scenarios typically operate under power constraints while guaranteeing maximum performance and efficiency. The NVIDIA Jetson series provides a good embedded and edge computing platform to evaluate AI model performance in such constrained environments.
@@
The experiments were conducted to assess the model’s inference speed on NVIDIA Jetson platforms, including Jetson Nano and Jetson AGX Orin, across power modes ranging from 5W to 60W. Each model was evaluated in full-precision (FP32) and mixed-precision modes. The evaluations were performed using the ImageNet-1K dataset with a batch size of 1 for the Jetson Nano and 128 for the Jetson AGX Orin. Both devices installed the latest NVIDIA JetPack SDK. The Jetson Nano used JetPack SDK 4.6.6 and the Jetson AGX Orin utilized JetPack SDK 6.1. Each experiment was repeated 100 times and the reported results represent the average of these repetitions to ensure statistical reliability.
Tables VIII and IX show the comparison of inference speed of different attention mechanisms obtained on ImageNet-1K using an NVIDIA Jetson AGX Orin/NVIDIA Jetson Nano GPU. In both FP32 and mixed precision, under the backbone of CSWin-T24181, ELFATT achieves the highest FPS. In mixed precision, ELFATT is significantly faster than all other attention mechanisms, as shown in Tables VIII and IX. Also, from Tables VIII and IX, FLatten does not have a significant acceleration effect compared to the vanilla softmax-based attention mechanism and is even slower than the vanilla softmax-based attention mechanism in mixed precision on the edge GPUs. Because FLatten uses a lot of normalization and reshaping operations in addition to linear attention. It further shows that these kinds of linear attention are not linear for all platforms. Even compared to EfficientViT-B2 [50], to achieve similar accuracy, ELFATT is significantly faster than EfficientViT-B2 in mixed precision on the edge GPUs. Fig. 4 shows the speed comparison of different attention mechanisms across various power modes from 5W to 60W by using NVIDIA Jetson Nano or NVIDIA Jetson AGX Orin. ELFATT consistently achieves 1.6x to 2.0x speedups compared to other attention mechanisms in various power modes from 5W to 60W. Compared to EfficientViT-B2, in mixed precision, ELFATT performs on par with EfficientViT-B2 in low power conditions, such as the 5W mode. As power increases, the speed growth rate of ELFATT surpasses that of EfficientViT-B2, especially in high-power modes like 50W and 60W, where it demonstrates a significant performance advantage. In mixed precision, ELFATT outperforms all other models in terms of speed across all tested power modes, establishing itself as the fastest option in both high-performance and power constrained scenarios.
V Conclusion
In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism for ViTs to achieve low memory I/O operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT is compatible with FlashAttention. Using FlashAttention-2 acceleration, ELFATT still offers 2-3x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT without using FlashAttention-2 acceleration is even faster than the vanilla softmax-based attention mechanism using FlashAttention-2 acceleration on both low-resolution and high-resolution vision datasets, which shows great potential of ELFATT. What’s more, on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W.

### v4

@@ Appendix D Ablation Study of Number of Levels Using the ELFATT Module
To validate the effect of the number of ELFATT modules used to replace the VaniATT modules at different levels of vision transformers, we compared the performance of CSWin-T using a different number of ELFATT modules at different levels. As shown in Table 11, with the number of levels using ELFATT modules increasing, the inference speed increases without using FlashAttention-2. The difference between CSWin-T using 3 levels of ELFATT modules and 4 levels of ELFATT modules is not significant. The fourth level of CSWin-T is composed of only one module and the sequence length is only 49 which is too short to achieve a significant acceleration effect. Another thing that can be found in Table 11 is that with an increase in the number of levels using ELFATT modules, CSWin-T using FlashAttention-2 achieves the fastest inference speed when using two levels of ELFATT modules. Because the sequence lengths of the first two levels are 3136 and 784, respectively, which are significantly longer than the last two levels (the $3^{\rm rd}$ level: 196, and the $4^{\rm th}$ level: 49). That is also the reason why some efficient attention mechanisms, such as Agent (Han et al., 2025) and FLatten (Han et al., 2023), only replace some of levels by their efficient attention modules. We also observed that with an increasing number of levels using ELFATT modules, the performance shows a gradual decline. To address this defect, we introduced a hybrid architecture in the third level which replaces half of the ELFATT modules at this level by VaniATT modules. The order of composition is as follows: VaniATT-ELFATT-...-VaniATT-ELFATT. The reason is that the third level of the vision transformer is usually much deeper than other levels and the sequence of this level is also much shorter than the first two levels. VaniATT at this level will not affect the speed too much and can help the model to converge faster. Swin-T-ELFATT, Swin-B-ELFATT, and CSWin-B-ELFATT also use a pure ELFATT architecture in the first two levels and a hybrid architecture in the third level. All variants of CSWin-T using ELFATT modules to replace VaniATT modules are faster than CSWin-T-GLOBAL.
Appendix E Ablation Study of Different $c_{1}$ and $c_{2}$
@@
Table 12 shows the performance comparison of CSWin-T-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. When $c_{1}=1\times c$ and $c_{2}=0\times c$, ELFATT becomes EFFATT, and when $c_{1}=0\times c$ and $c_{2}=1\times c$, ELFATT becomes a local window-based attention mechanism which can be regarded as a simpler version of the local window-based attention mechanism used in Swin. As shown in Table 12, $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ achieve better performance than other combinations of $c_{1}$ and $c_{2}$. Without using FlashAttention-2, the difference between $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ in terms of speed is not significant. Using FlashAttention-2, $c_{1}=0\times c$ and $c_{2}=1\times c$, and $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$ achieve close speed and are significantly faster than other combinations.
Appendix F Ablation Study of Different Patch Partition/Merging Methods
Table 13 shows the performance comparison of Swin-T-ELFATT using different patch partition/merging methods on ImageNet-1K. “concat” (Liu et al., 2021) denotes the concatenation-based patch partition/merging method of Swin. “conv” (Dong et al., 2022) denotes the convolution-based patch partition/merging method of CSWin. It can be seen that “conv” significantly outperforms “concat” in terms of speed and accuracy. Hence, in this paper, all ELFATT variants use “conv” for patch partition/merging.
Appendix G Speed Comparison on Edge GPUs: Experiment Settings
@@
In addition to high-performance computing applications, model deployment should also be evaluated in emerging edge scenarios such as robotic vision, unmanned aerial vehicles (UAVs), and autonomous driving. These scenarios typically operate under strict power constraints while striving to maximize performance and efficiency. The NVIDIA Jetson series provide a good embedded and edge computing platform to evaluate AI model performance in such resource-constrained environments.
Table 13. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT using different patch partition/merging methods in Swin-T on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on 1 NVIDIA H20 (96 GB) GPU.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
concat $224^{2}$ 82.3 2027/2769 imgs/s 28M 6.81G/4.42G
conv $224^{2}$ 82.7 2884/3159 imgs/s 30M 4.99G/4.67G
G.1. Experiment Settings
@@
The experiments were conducted to evaluate the inference speed of the model on NVIDIA Jetson platforms, specifically Jetson Nano and Jetson AGX Orin, across power modes ranging from 5W to 60W. Each model was evaluated in full precision (FP32) and mixed precision modes. The evaluations were performed on the ImageNet-1K dataset with batch sizes of 1 for Jetson Nano and 128 for Jetson AGX Orin. Both devices installed the latest NVIDIA JetPack SDK (The Jetson Nano used JetPack SDK 4.6.6 and the Jetson AGX Orin utilized JetPack SDK 6.1). Each experiment was repeated 100 times and the reported results represent the average values to ensure statistical reliability.
Figure 5. The speed comparison of different attention mechanisms across power modes from 5W to 60W on NVIDIA Jetson Nano and AGX Orin under power-constrained inference scenarios.
Appendix H Stable Diffusion Acceleration Using ELFATT: Experiment Settings and Ablation Studies
We also introduced ELFATT to the acceleration of stable diffusion (SD) (Rombach et al., 2022). The original EFFATT-based global linear attention head and LePE of ELFATT cannot be used without training. We used Agent (Han et al., 2025) to replace the original EFFATT-based global linear attention head. The reason is that: (1) Agent is an improved version of EFFATT, and it only introduces a small agent matrix as the auxiliary key or query matrix to be multiplied with the full query or full key matrix to reduce its dimenison before the softmax normalization; (2) Agent can be used to replace the self-attention block of SD without training. Hence, Agent can be introduced to replace the original EFFATT-based global linear attention head of ELFATT. LePE is removed. Finally, in the self-attention block of SD, Eq. (10) of the main body of the paper will become as follows,
(18) $\begin{split}{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\approx&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\mathcal{A}}^{\top}\right){\rm exp}\left(\bar{\mathcal{A}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$
Eq. (11) of the main body of the paper will become as follows,
(19) $\begin{split}\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\approx&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\mathcal{A}}^{\top}\right){\rm exp}\left(\bar{\mathcal{A}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$
where $\bar{\mathcal{A}}={\rm Pooling}\left(\bar{\textit{{Q}}}\right)\in\mathbb{R}^{m_{1}\times c_{1}}$, ${\rm Pooling}(\cdot)$ denotes the average pooling function, and $m_{1}\ll m$.
Table 14. The comparison of stable diffusion (SD) v1.5 (Diffusers v0.32.1) using ELFATT with different block sizes in FP16 precision where merging ratio is 0.5. Memory consumption for each variant indicates how much GPU memory increases when the batch size is raised by 1. All variants are accelerated by FlashAttention-2. The processing time of each variant is obtained using a batch size of 48 on 1 NVIDIA Tesla A100 (40 GB) GPU.
Block size FID $\downarrow$ Time (s/img) Memory (GB/img)
32 30.8 0.710 0.65
56 30.7 0.709 0.65
81 30.5 0.708 0.65
128 30.4 0.707 0.70
227 30.5 0.705 0.70
512 30.5 0.709 0.70
H.1. Experiment Settings
The experiments were carried out to assess the performance and acceleration effect of ELFATT in a stable diffusion task using 1 NVIDIA Tesla A100 (40 GB) GPU. Following the experiment settings of (Han et al., 2025) and (Bolya and Hoffman, 2023), we applied ELFATT (Eqs. (18) and (19)) on stable diffusion v1.5 using the pipeline of Diffusers v0.32.1 and called the proposed stable diffusion acceleration method ELFATT-SD. We compared ELFATT with Agent (Agent-SD) (Han et al., 2025) and ToMe (ToMe-SD) (Bolya and Hoffman, 2023). We used ImageNet-1K labels as prompts to generate 2000 images with a resolution of $512^{2}$ (2 images per class). The inference steps (50) (Liu et al., 2022b) and the cfg scale (7.5) (Dhariwal and Nichol, 2021) remain the same as Agent and ToMe. Similar to Agent-SD, ELFATT-SD is also developed based on the token merging method ToMe-SD. We also used the hybrid architecture of Agent-SD which means ELFATT was used in early inference steps, and after that vanilla ToMe was applied for the remaining inference steps. The ratio of inference steps used for ELFATT is 20%. The evaluation metric used is the FID score (Heusel et al., 2017) which is used to evaluate the difference between 2000 generated images and 50000 images of the ImageNet-1K validation set.
Figure 6. The visual comparison of images generated by Agent-SD (merging ratio: 0.5), ELFATT-SD (merging ratio: 0.5), SD, and ToMe-SD (merging ratio: 0.5).
Table 15. The comparison of SD using ELFATT in different inference steps with BF16 precision (merging ratio is 0.5).
Steps FID $\downarrow$ Time (s/img) Memory (GB/img)
0 (ToMe-SD) 30.9 0.714 0.70
10 30.4 0.707 0.70
20 30.6 0.702 0.60
H.2. Ablation Study of Different Block Sizes
Table 14 shows the effect of different block sizes used in ELFATT on the performance of the stable diffusion task. When the block size increases, the speed is first improved and then declined. When the block size is 227, the speed is the fastest. The speed of the block size 128 is the second fastest and this block size has the lowest FID score. It seems that when the block size becomes smaller than 227, the inference speed is slowed, which may be caused by FlashAttention-2 and GPU architectures. In summary, the speed difference of the block size range from 32 to 512 is not significant. The reason is that: (1) Even the largest block size 512 is 8x smaller than the original sequence length 4096. (2) Only the first 20% inference steps were applied to use ELFATT for acceleration. However, our method is still significantly faster than other methods, especially the baseline which shows the effectiveness of ELFATT for acceleration of SD.
H.3. Ablation Study of Number of Inference Steps Using ELFATT
Table 15 shows the effect of different inference steps using ELFATT on the performance of the stable diffusion task. When 0 inference steps using ELFATT are adopted, ELFATT-SD becomes ToMe-SD. With more inference steps using ELFATT being adopted, the speed becomes faster; however, the FID score is reduced first and then increased. When early 10 (20%) inference steps using ELFATT are adopted, the FID score is the lowest.